A body rolled on a surface with a speed of 10m/s comes to a rest after covering a distance of 50 m . Find the coefficient of friction Take g=10m/s^2

Grade:11

2 Answers

Abhishek singh
6 years ago
Here friction produce retardation which is úg From question U=10m\sV=0S=50ma=-úgPutting these value in equation of motionV^2= U^2+2asWe getú=0.1
muneeb
6 years ago
u=10m/s. s=50. v=0. 
then by third equation of motion 
v2=u2+2as
after solving we get 
a=-1m/s2
then u*=f/R=ma/mg=.1m/s  (g=10m/s2)